# Emergence of topological states in relaxation dynamics of interacting bosons

### v1

@@ Acknowledgements
We acknowledge useful discussions with Zhen-Sheng Yuan, Xiong-Jun Liu, Bing Yang, and Jun-Jun Xu. This work is supported by the National Natural Science Foundation of China under Grants Nos. 12374252, 12074431, and 12174130, and the Excellent Youth Foundation of Hunan Scientific Committee under Grant No. 2021JJ10044. Numerical calculations were performed on the TianHe-1A cluster of the National Supercomputer Center at Tianjin.
Data availability
The authors declare that the data supporting the findings of this study is available within the article and its Supplementary Information. Extra data is available from the corresponding authors upon reasonable request.
Author Contribution
Y.L. conceived the project. W. H., X.-C. Y., and R.C. performed the calculations with input from others. All authors contributed to analysis of the results and writing of the paper.
Competing interests
The authors declare no competing interests.
Supplementary Information
@@
In this Supplementary Information, we provide more details about the ground-state properties, and far-from-equilibrium many-body dynamics, and construct effective models to provide additional insight into the physics. Without loss of generality and unless otherwise specified, we focus on the two cases studied in the main text, i.e., $J_{1}/J_{2}=0.2$ and $U/J_{2}=8$ in the nontrivial regime, and $J_{1}/J_{2}=0.8$ and $U/J_{2}=6$ in the trivial regime.
@@ II.3 Dynamics of parity and density-wave correlations
II.4 Dynamics of the entanglement entropy
@@
II.5 Dynamics of the entanglement spectrum

### v2

@@ Abstract
We acknowledge useful discussions with Zhen-Sheng Yuan, Xiong-Jun Liu, Bing Yang, and Jun-Jun Xu. This work is supported by the National Natural Science Foundation of China under ation of China under Grants Nos. 12374252, 12074431, and 12174130, and the Excellent Youth Foundation of Hunan Scientific Committee under Grant No. 2021JJ10044. Numerical calculations were performed on the TianHe-1A cluster of the National Supercomputer Center at Tianjin.
Supplemental Material
@@
In this Supplemental Material, we provide more details about the ground-state properties, and far-from-equilibrium many-body dynamics, and construct effective models to provide additional insight into the physics. Without loss of generality and unless otherwise specified, we focus on the two cases studied in the main text, i.e., $J_{1}/J_{2}=0.2$ and $U/J_{2}=8$ in the nontrivial regime, and $J_{1}/J_{2}=0.8$ and $U/J_{2}=6$ in the trivial regime.
@@ II.3 Dynamics of parity and density-wave correlations
II.4 Initial-state dependence of relaxation dynamics
As is widely recognized, initial states typically influence many-body dynamics in time evolution. It is pertinent to examine the impact of the initial state on string, parity and density-wave correlations. With the performance of numerical calculations in different initial states, the long-range correlations remain robust to the initial states once singly occupied on the two sites adjoined by $J_{2}$-bonds. As shown in Fig. S6, the string correlations of the long-time evolved stationary states exhibit nonzero values for different lengths, while parity and density-wave orders decay to zero quickly. Note here that these initial states satisfy half filling in the bulk naturally, reminiscent of ground-state topology. The filling at the edges is not essential to this bulk phenomenon, which is attributed to the distinct relaxation times of the bulk and boundary states. The robustness of initial states is beneficial to the experimental research for observing the exotic dynamics.
Figure S6: Dynamics of the string, parity, and density-wave correlations. From top to bottom, time evolution of the three quantities for four different lengths $|i-j|$ in the nontrivial regime. Parameters: $L=20$, $U/J_{2}=8$, and $J_{1}/J_{2}=0.2$. The results are obtained with bond dimension up to $1000$.
We have performed additional calculations to map out the phase diagram for different initial states. It is found that the phase diagram is remarkably robust in the timescale $t=100/J_{2}$, which is a typical value accessible in ultracold atom experiments. Numerical results about the string order are shown in Fig. S7. We have checked many different values of $J_{1}/J_{2}$ with $U/J_{2}=8$. All initial states are chosen to satisfy the constraint mentioned above: there is only one boson in every two sites connected by a $J_{2}$ hopping in the bulk. The phase boundary is almost independent of the initial state.
Figure S7: String order as a function of the hopping strengths for different initial states. String order of the long-time stationary state at $t=100/J_{2}$. The interaction strength is fixed at $U/J_{2}=8$ and the hopping strength $J_{1}/J_{2}$ is varied. The solid lines correspond to the $L=20$ system with multiple initial states given in the inset and the dashed line corresponds to the $L=32$ system with initial state $\left|1010\cdots\right\rangle$. The MPS bond dimension is 1000.
II.5 Dynamics of the entanglement entropy
@@
II.6 Dynamics of the entanglement spectrum